# Modelling, Design Optimization and Prototype development of Knee Exoskeleton

### v1

@@ 6 Graphical Simulation
In optimization, particularly in iterative approach, the global solution may not necessarily the correct solution. The global minima might not be feasible in the context of design of the exoskeleton.
Figure 7: Graphical Simulation of optimal exoskeleton at $\theta$ = $2^{\circ}$
Figure 8: Graphical Simulation of optimal exoskeleton at $\theta$ = $70^{\circ}$
Figure 9: Graphical Simulation of optimal exoskeleton at $\theta$ = $110^{\circ}$
Figure 10: Graphical Simulation of optimal exoskeleton at $\theta$ = $148^{\circ}$
In optimization, particularly in iterative approach, the global solution may not necessarily the correct solution. The global minima might not be feasible in the context of design of the exoskeleton. The solution obtain is an approximate solution which is local minima as shown in the figure 5. To verify the practicality of the solution, author constructed a method in MATLAB to demonstrate the geometrical simulation based on optimal dimension which is shown in figure 7. In this figure, the ordinate and abscissa represents the dimensions of the links. The red circle shows the knee joint, around which the rotation will occur. In figure 7, there is 2 degree of rotation i.e. the human is almost standing. As the human start siting, the angle of rotation ($\theta$) increases as you can see in figure 8 and 9. The figure 10 shows the maximum angle achieved by the exoskeleton ( $\theta$ = $148^{\circ}$ ) before reaching the singularity position. The rotation of exoskeleton is actuated by linear actuator as shown in figure 7 to 10. The thick red line represent the cylinder of the linear actuator and black line represents piston. The combine length of piston and cylinder is represented as d in the figures. The figure 10 validate that the optimal design of exoskeleton with same stroke length (d) gives more range of rotation than previous design.
Figure 11: Simulation with mannequin
@@ 7 Gait Analysis
In this section, there is analysis of Sit to Stand motion of human being and LAKE. The Gait analysis is crucial while designing the exoskeleton.Understanding the gait mechanics is particularly important for designing exoskeletons used in rehabilitation. Gait analysis helps identify abnormal movement patterns and can guide the design of exoskeletons to correct these patterns, aiding in the recovery process. Additionally, it helps prevent further injuries by ensuring that the exoskeleton does not impose unnatural stresses on the body. Here by analyzing the individual gait pattern, the exoskeleton can be constructed with minimum misalignment.
To perform the gait analysis of human being for Sit to Stand motion, The author performed experiment in which subject is performing sit to stand motion. This motion has captured with the help of camera. Figure 13 illustrates the setup where the sit-to-stand motion was captured using a high-resolution camera. A reflective marker was strategically placed on the knee joint of the subject to ensure precise tracking. The TRACKER software processed the video footage frame by frame, extracting the trajectory data of the marker over time. The markers trajectory shown in figure are the trajectories of the joint. To analyse the motion of these coordinates, the angles has to be calculated.
Figure 12: Vector Allocation
Figure 13: Sit to Stand motion of human
This was done by defining vectors between key anatomical landmarks: $\vec{AK}$ (ankle to knee), $\vec{KH}$ (knee to hip), and $\vec{HW}$ (hip to waist) as shown in fig 12. These vectors were used to determine the angular displacements and provide a comprehensive view of the joint mechanics. The Knee joint angle is calculated by formula
@@
where, $\theta$ is knee joint angle. This thorough and systematic approach offers significant insights into the bio-mechanical interactions between human joints and exoskeleton systems, emphasizing the importance of precise alignment in enhancing the efficacy and safety of exoskeletons in practical applications.
@@
Similarly, same subject has performed the sit to stand motion with exoskeleton. This exoskeleton is dummy model made up from cardboard. Its sole purpose is to get the trajectory of optimized dimension based exoskeleton. In figure 14, subject is performing sit to stand motion with exoskeleton. The red mark on exoskeleton is exoskeleton’s knee joint. Here, it is an assumption that red mark act as knee joint. The actual knee joint can be instantaneous center of four bar linkage. It is because the instantaneous centre of four bar linkage follow the same trajectory as human’s knee joint during sit to stand motion.
The data obtained from the experiment is used to draw the trajectory of the human’s knee joint and exoskeleton’s knee joint. The concept behind these generation of trajectories are to get some idea on misalignment between them. The yellow line trajectory shows the motion of knee joint during sit to stand motion as shown in figure 15. The orange line shows the trajectory of the exoskeleton’s knee joint. The figure 16 shows the relative error between both trajectories. Since, both trajectory are similar, the relative error is zero at most position. This similarity of trajectory shows that the better understanding of misalignment between them. This similarity also conclude that the misalignment is quite less between the subject and exoskeleton. The sudden spike in the error as shown in figure 16 is due to the human error during the experiment. During experiment, sudden changes in movement of subject or friction between the links can cause a slight deviation of position of exoskeleton’s knee joint.
Figure 15: Motion analysis comparison
Figure 16: Relative error between both trajectory
@@ 8 Prototype Development
The prototype of Linear Actuator based Knee joint Exoskeleton (LAKE) has been developed. In this work, the micro-controller used is Arduino UNO as shown in figure 17. The Arduino UNO is an open-source electronics platform based on easy-to-use hardware and software.The primary purpose of the Arduino UNO is to facilitate the development of digital devices and interactive objects that can sense and control objects in the physical world.
@@
This Arduino works on 5V power supply. The power supply has been provided by Laptop. To control the Linear actuator, the motor driver L298N has been used. Motor drivers are crucial components in the field of electronics, particularly in robotics and automation. They serve as the intermediary between microcontrollers (such as Arduino boards) and motors, enabling the control of motor direction and speed. This is important because microcontrollers often cannot supply the necessary current to drive motors directly. Motor drivers amplify the control signals from the microcontroller to provide the appropriate power to the motors.The author used L298N motor driver because of its high current capability, as the L298N can drive motors with up to 2A of current per channel, accommodating a broad range of motor types and sizes. The module operates over a wide voltage range, handling motor supply voltages from 5V to 35V, which adds to its adaptability in various projects.
Speed control is another critical feature of the L298N, achieved through pulse-width modulation (PWM). By varying the duty cycle of the PWM signal applied to the enable pins (ENA and ENB), users can finely adjust motor speeds. The module also includes thermal protection to prevent damage from overheating, a crucial aspect for maintaining reliability and longevity in demanding applications. Additionally, the L298N is designed for easy integration with straightforward onboard terminals and connectors, simplifying connections to motors and power supplies. The power supply is given from adaptor which convert 220V AC into 12V 2A DC, which is suitable output for our motor driver.
@@
After assembling all the components shown in figure 17, the C based code has been written in Arduino IDE. This code is uploaded to the Arduino which drives the motor driver and further the exoskeleton as shown in figure LABEL:sts_exo. The exoskeleton wear by subject in figure LABEL:sts_exo is working at 5mm/s stroke velocity. There is no support except exoskeleton has been taken by subject. The forces generated by the actuation of actuator pushes the subject upward. Furthermore, this prototype has been used to validate the graphical simulation as shown previously in figure LABEL:fig12. The angle measured at given stroke length (252mm) as shown in figure 19 and the angle achieved during Graphical simulation by providing same stroke length (252mm) comes similar which verify our previous simulation and analysis.
@@ 9 Conclusion
The mechanical design of existing knee joint exoskeleton has been optimized for maximum range of motion during Sit to Stand motion. The optimization technique used is Interior Point method. The ranges used in this method can be change according to the measurements of the subject. Furthermore, the results have been shown by constructing the graphical simulation and simulation with mannequin. Additionally, The prototype has been developed on the basis of the optimization and simulations which verify the maximum range of motion of knee exoskeleton.

### v3

@@ 5 Graphical Simulation
In constrained optimization problem, the global solution may not necessarily be the desired solution. The global minima might not be feasible in the context of design of the exoskeleton. Therefore, authors have constructed an algorithm to check the feasibility of the solution by providing a geometrical simulation of the knee exoskeleton based on optimal dimension as shown in Fig 5. The obtained feasible solution based on the algorithm, is an approximate solution which is a local minima, as shown in the Fig 4(a).
(a) $\theta$ = $2^{\circ}$
(b) $\theta$ = $70^{\circ}$
(c) $\theta$ = $110^{\circ}$
(d) $\theta$ = $148^{\circ}$
Figure 5: Graphical Simulation of optimal exoskeleton at (a) $\theta$ = $2^{\circ}$ (b) $\theta$ = $70^{\circ}$ (c) $\theta$ = $110^{\circ}$ (d) $\theta$ = $148^{\circ}$
@@ 6 Gait Analysis
This section provides the misalignment study of proposed Linear Actuator based Knee Exoskeleton (LAKE) for sit to stand (STS) motion. This can be done by analyzing the human knee’s STS motion and LAKE’s STS motion. To perform the gait analysis of human being for STS motion, an experiment is conducted in which subject is performed sit to stand motion. This motion has captured with the help of a high-resolution camera as shown in Fig 6(b). A reflective marker is strategically placed on the knee joint of the subject to ensure precise tracking. The TRACKER software processed the video footage frame by frame, extracting the trajectory data of the marker over time. To analyze the motion of these coordinates, the angles has to be calculated.
(a) Vector Allocation
(b) Experiment with human (alone) during STS motion
Figure 6: Optimum dimension based knee exoskeleton
This was done by defining vectors between key anatomical landmarks: $\vec{AK}$ (ankle to knee), $\vec{KH}$ (knee to hip), and $\vec{HW}$ (hip to waist) as shown in Fig 6(a). These vectors were used to determine the angular displacements and provide a comprehensive view of the joint mechanics. The knee joint angle is calculated as
@@
where, $\theta$ is the knee joint angle.
@@
Similarly, same subject has performed the STS motion with exoskeleton. This exoskeleton is a dummy model made up from wooden cardboard. Its purpose is to get the trajectory of optimized dimension based knee exoskeleton. In the Fig 7, subject is performing sit to stand motion with exoskeleton. The data obtained from the experiment is used to draw the trajectory of the human’s knee joint and exoskeleton’s knee joint. The concept behind these generation of trajectories are to get some idea on misalignment between them. The yellow line trajectory shows the motion of knee joint during STS motion as shown in Fig 8(b). The orange line shows the trajectory of the exoskeleton’s knee joint. The Figure 8(a) shows the relative error between both trajectories. Since, both trajectory are similar, the relative error is found as zero at most of the positions. This similarity conclude that the misalignment is minimum between the subject and exoskeleton. The sudden spike in the error as shown in Fig 8(a) is due to the human error during the experiment.
(a) Relative error between both trajectory
(b) Comparison of the STS motion with and without exoskeleton
Figure 8: Gait Analysis (a) Relative error between two trajectories (b) Trajectory of human knee adn exoskeleton knee joint.
@@ 7 Prototype Development
The prototype of Linear Actuator based Knee joint Exoskeleton (LAKE) has been developed. In this work, the micro-controller used is Arduino UNO, as shown in Fig 9.
@@
The power supply has been provided by laptop. To control the linear actuator, the motor driver L298N has been used. This is important because microcontrollers often cannot supply the necessary current to drive motors directly. Motor drivers amplify the control signals from the microcontroller to provide the appropriate power to the motors.The author used L298N motor driver because of its high current capability, as the L298N can drive motors with up to 2A of current per channel, accommodating a broad range of motor types and sizes. The module operates over a wide voltage range, handling motor supply voltages from 5V to 35V, which adds to its adaptability in various projects. Speed control is another critical feature of the L298N, achieved through pulse-width modulation (PWM). By varying the duty cycle of the PWM signal applied to the enable pins (ENA and ENB), users can finely adjust motor speeds.
@@
After assembling all the components shown in Fig 9, the C language based code is written in Arduino IDE. This code is uploaded to the Arduino which drives the motor driver and further the exoskeleton as shown in Fig 10. The exoskeleton worn by the subject in Fig 10 operates at a stroke velocity of 5mm/s5\;mm/s. The subject receives no external support other than the exoskeleton itself. The upward motion is facilitated solely by the forces generated through actuator actuation. Furthermore, this prototype has been used to validate the graphical simulation as shown in Fig 5 . The angle achieved by the LAKE in both the graphical simulation and the physical prototype is found to be equal at the given stroke length, as shown in the Fig 11.
@@ 8 Conclusion
In this paper, the optimization algorithm has been constructed which optimize the design of closed loop 4-bar mechanism knee exoskeleton for maximum range of rotation. To achieve that, the algorithm used the nonlinear equation derived from kinematic modeling of exoskeleton and optimized it using interior point method and then check the feasibility of result by simultaneously constructing the geometrical simulation of the optimized knee exoskeleton for maximum range of rotation. This optimized exoskeleton has been modeled using CAD software to further validate its maximum range of motion and feasibility by harnessing on the human mannequin based on average human measurement. To investigate joint misalignment, an experiment was conducted in two stages: first, with a human knee joint performing the sit-to-stand (STS) motion, and then with a dummy fitted with the optimized knee exoskeleton. The marker on the knee joint and software provide us the trajectory of the human knee joint and knee exoskeleton joint. The relative error between the two trajectories ensured that knee exoskeleton joint and human joint are almost in sync during STS motion. Furthermore, to validate the optimized results, a prototype of Linear actuator based Knee exoskeleton (LAKE) was developed. The prototype utilizes an Arduino microcontroller and operates under an open-loop control scheme to perform the STS motion.